# BadPart: Unified Black-box Adversarial Patch Attacks against Pixel-wise Regression Tasks

### v1

@@ Abstract
Adversarial Robustness, Patch Attack, Black-Box Attack, Monocular Depth Estimation
@@ 1 Introduction
Previous adversarial patch attacks on pixel-wise regression models (Cheng et al., 2022) have primarily been conducted in a white-box setting, where the attacker has complete knowledge of the target model and can use gradient-based optimization to create the patch and compromise the model. However, this white-box assumption is not always feasible. For instance, models used in autonomous driving systems are typically closed-source (Tesla, ). Moreover, some models are provided as an online service, preventing users from accessing their internals model structures (Google3DPortrait, ). In typical scenarios, black-box attacks are more realistic and could pose a greater threat to the security of these tasks. Existing black-box attack techniques are either based on training a substitute model or iteratively querying the model. In this study, we focus on the query-based approach as it reduces the cost of model training and only assumes that the attacker has access to the model output. This leads us to a pivotal inquiry: “can we craft an adversarial patch to compromise the pixel-wise regression model through iterative querying in a black-box manner?
@@
To address this issue and enhance the scalability of black-box patch generation, we leverage the domain knowledge of pixel-wise regression tasks and propose a novel square-based adversarial patch optimization framework. Specifically, in each iteration, we consider a small square area within the patch region to reduce the potential search space, then we introduce a batch of random noises on the square area to estimate the gradients of the region by evaluating the score of each noise in terms of attack performance. We sample the location of the square area probabilistically based on the pixel-wise error distribution, and propose novel score adjustment procedures for more precise gradient estimation. The source code is available. Figure 1 presents our attack performance. In summary, our contributions are as follows:
@@ 2 Background and Related Work
Pixel-wise Regression Tasks. Pixel-wise regression tasks generate continuous values for each input pixel, differing from pixel-wise classification tasks like semantic segmentation (Yu et al., 2018; Zhou et al., 2022), which assign a discrete class label to every pixel. Representative tasks of this type include monocular depth estimation (MDE) (Moon et al., 2019; Watson et al., 2019; Wang et al., 2023a), optical flow estimation (OFE) (Teed & Deng, 2020; Ilg et al., 2017; Dosovitskiy et al., 2015), and surface normal estimation (SNE) (Zeng et al., 2019; Lenssen et al., 2020; Bae et al., 2021). In MDE models, the output comprises the estimated pixel-wise distance between the 3D scenario and the camera capturing the input image, with each pixel corresponding to a distance estimation. OFE models use two consecutive image frames as input and output the estimated motion of pixels (the “optical flow”) between the two frames. For each pixel in the first frame, a 2D vector is estimated, indicating its offset to the corresponding pixel in the second frame. SNE models output the estimated orientations of the surfaces in the input image, described with the ”normal vectors” in 3D space that are perpendicular to the surface at the locations of pixels. Given that the first two tasks (MDE and OFE) have broader and more security-critical applications, such as autonomous driving (Karpathy, 2020), visual SLAM (Wimbauer et al., 2021), video composition (Liew et al., 2023), and augmented reality (Bang et al., 2017), our discussion primarily focuses on these two tasks. However, our proposed attack is a unified approach and can be readily applied to other pixel-wise regression tasks like SNE.
@@ 3 Problem Formalization
$\mathcal{F}(\mathbf{D})[i,j]$ $=Mean(\{\mathbf{\mathbf{D}}[k,1,i,j]~{}|~{}k=1...n\})$ (8)
$\mathcal{F}(\mathbf{D})[i,j]$ $=Mean(\{\|\mathbf{D}[k,:,i,j]\|_{2}~{}|~{}k=1...n\})$ (9)
@@ 4.2 Probabilistic Square Sampling
The sampling algorithm is designed to enhance the probability of selecting locations within the patch region that are more vulnerable to adversarial perturbations. As indicated in (Cheng et al., 2022) and (Cheng et al., 2023), the identification of vulnerable areas on images is critical for improving attack performance. To ascertain these vulnerable areas in the beginning, we employ an initialization phase ($K$ iterations), wherein we randomly sample the square location in a uniform manner. This period is denoted as $\mathbf{p^{(1)}}$ in Figure 2. After the initialization phase, the pixel-wise error map $\mathbf{M}$ caused by the latest patch is utilized as an indication of the vulnerable areas and we leverage this map to calculate the probability distribution of location sampling. Those areas with larger errors obtain higher probability. The background color of the patch $\mathbf{p^{(1)}},\mathbf{p^{(2)}}$ and $\mathbf{p^{(3)}}$ in Figure 2 illustrate the sampling probability distribution of square locations.
@@ 4.3 Score-based Gradient Estimation
15:     $[\Delta]^{+}\leftarrow[\Delta]^{+}/Max([\Delta]^{+})$;   $[\Delta]^{-}\leftarrow[\Delta]^{-}/|Min([\Delta]^{-})|$
@@
5 Evaluation
In this section, we evaluate BadPart on 2 kinds of tasks including 7 subject models. We compare with 3 baseline black-box attack methods and a white-box one. A set of ablation studies are discussed and the source code is provided.
5.1 Experiment Setup
@@ 57 $\times$ 57 Patch (1%) 80 $\times$ 80 Patch (2%) 100 $\times$ 100 Patch (3%)
Models GA HB S-RS Ours WB GA HB S-RS Ours WB GA HB S-RS Ours WB
@@
  • * Bold texts denote the best attack performance among black-box methods. Abbreviations. GA: GenAttack (Alzantot et al., 2019), HB: HardBeat (Tao et al., 2023), S-RS: Sparse-RS (Croce et al., 2022), WB: White-box Attack for reference.
@@
Tasks & Models. We evaluate BadPart on two pixel-wise regression tasks of MDE and OFE. For MDE, we use Monodepth2 (Godard et al., 2019), DepthHints (Watson et al., 2019), SQLDepth (Wang et al., 2023b) and PlaneDepth (Wang et al., 2023a) as the target models. For OFE, we attack FlowNetC (Dosovitskiy et al., 2015), FlowNet2 (Ilg et al., 2017) and PWC-Net (Sun et al., 2018). These models are selected since they cover both the popular and SOTA models, and we attack the publicly available models with the highest input resolutions from their official repositories. MDE models were trained on the KITTI depth prediction dataset (Uhrig et al., 2017) and OFE models were trained on the KITTI flow 2015 (Menze et al., 2015).
@@
Baselines & Metrics. There are no direct baselines available due to a lack of previous research on black-box attacks against pixel-wise regression tasks. Hence we adapt three black-box patch attacks on image classification to our scenario as baselines. These include Sparse-RS (Croce et al., 2022), a soft-label attack that employs random search and is akin to the patch-attack variant of SquareAttack (Andriushchenko et al., 2020); HardBeat (Tao et al., 2023), a SOTA hard-label attack that estimates gradients; and GenAttack (Alzantot et al., 2019), a conventional approach using genetic algorithms. We also compare BadPart with the white-box attack as reference. We employ the mean depth estimation error (DEE, unit: meters) for MDE models and the mean end point error (EPE, unit: pixels) for OFE models as metrics to denote attack performance (the higher the better), which are aligned with the metrics used in prior white-box attacks on the two tasks (Cheng et al., 2022; Ranjan et al., 2019). They are calculated on the patched area and reflect the error caused by the generated adversarial noise.
@@ 5.2 Attack Performance
We compare the attack performance of BadPart with other baseline methods. We report the maximum attack performance of each method until convergence or after 1000K queries, whichever first. We also include the performance of white-box attacks as references. Three different patch sizes are evaluated and the patch locates at the center of the image. Table 1 reports the result. The first column denotes various pixel-wise regression models under attack, and the following columns represent attack performance of different methods and the white-box reference. As shown, BadPart obtains the best attack performance on all models under various patch sizes. The performance of BadPart is even close to the white-box attack reference on some models (e.g., Monodepth2, DepthHints and SQLDepth). Sparse-RS has the second best attack performance while GenAttack performs the worse and has nearly no effect. Figure 3 presents qualitative examples of the attack performance of different methods on PlaneDepth and FlowNetC. As shown, the first column denotes the benign scene and the MDE and OFE output. The following columns show the $2\%$ adversarial patches generated by BadPart and baseline methods, as well as the model output. In the first row, when the patch generated by BadPart is applied, the depth estimation of the patched area is significantly further than the actual distance (darker color denotes further distance estimation). In comparison, the patches generated by other methods cause less impact. In the second row, patches generated by BadPart and Sparse-RS have degraded the OFE performance significantly, making the result unusable. More qualitative and quantitative results are in Appendix C.
@@ 5.5 Attack Online Service
To evaluate the feasibility of BadPart in real-world applications, we conduct attacks on the Google API for portrait depth estimation (Google3DPortrait, ). Note that the model is not deployed by authors and we only query it for depth estimation. We utilize ten $256\times 192$ portrait images as training set and generate a $31\times 31$ patch using 50K queries to the API. Our attack goal is to minimize the estimated distance of the patched area on the portrait. BadPart has successfully reduced the mean depth estimation of the patched area by 43.5% from 0.431 cm to 0.243 cm. Figure 7 shows qualitative results. This adversarial example is available in the supplementary material, which can be uploaded to (Google3DPortrait, ) for efficacy validation. The adversarial patch can also be employed for beneficial purposes, such as privacy protection. By attaching the patch to personal images before publishing, individuals can prevent unauthorized use of their photos in such video composition services.
5.6 Defensive Discussion
As pioneers in the exploration of black-box adversarial patch attacks against pixel-wise regression tasks, we find ourselves in uncharted territory with no existing defense techniques specifically tailored for this context. Nevertheless, there are defense strategies designed against black-box attacks on classification models. For example, certifiable defenses such as PatchCleanser (Xiang et al., 2022) employ a mask to traverse all input positions, monitoring output class mutations to identify the most suspicious position. Furthermore, universal adversarial patch detection methods (e.g., SentiNet (Chou et al., 2020)) depend on the input features responsible for the predicted class to locate the patch. However, their reliance on class output renders them unsuitable for direct application to the pixel-wise regression tasks. In contrast, query-based defense techniques, designed to detect malicious queries by black-box attacks, may be more applicable to our context. Blacklight (Li et al., 2022), a leading defense of this type, leverages the similarity among different query inputs to detect black-box attacks. Its primary strategy involves calculating the hash representation of each incoming query and identifying an adversarial query if the hash matches any previous one. Blacklight’s efficacy is contingent on the similarity between two images in consecutive queries, which is a major feature of single-image black-box attacks (e.g., SquareAttack (Andriushchenko et al., 2020)). However, BadPart is a universal adversarial patch attack that does not depend on sample similarity, and the randomness in different samples could potentially enhance the universal effectiveness of the generated patch. Hence we add random noise on each attack sample to by-pass the defense. Additionally, the high resolution of our input images further diminishes the efficiency and efficacy of such a defense. To assess the defensive performance of Blacklight on BadPart, we have incorporated it in our framework and evaluated its detection rate for varying amounts of queries. The results are presented in Table 3. As shown, for both MDE and OFE tasks, the detection rate remains zero under 800K queries, while the attack performance is not affected and continues to increase with more queries. We discuss the limitations and our future work in Appendix D.

### v3

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Previous adversarial patch attacks on pixel-wise regression models (Cheng et al., 2022, 2023) have primarily been conducted in a white-box setting, where the attacker has complete knowledge of the target model and can use gradient-based optimization to create the patch and compromise the model. However, this white-box assumption is not always feasible. For instance, models used in autonomous vehicles are typically closed-source (Tesla, ). Moreover, some models are provided as an online service, preventing users from accessing their internals model structures (Google3DPortrait, ). In typical scenarios, black-box attacks are more realistic and could pose a greater threat to the security of these tasks. Existing black-box attack techniques are either based on training a substitute model or iteratively querying the model. In this study, we focus on the query-based approach as it reduces the cost of model training and only assumes that the attacker has access to the model output. This leads us to a pivotal inquiry: “can we craft an adversarial patch to compromise the pixel-wise regression model through iterative querying in a black-box manner?
@@
To address this issue and enhance the scalability of black-box patch generation, we leverage the domain knowledge of pixel-wise regression tasks and propose a novel square-based adversarial patch optimization framework. Specifically, in each iteration, we consider a small square area within the patch region to reduce the potential search space, then we introduce a batch of random noises on the square area to estimate the gradients of the region by evaluating the score of each noise in terms of attack performance. We sample the location of the square area probabilistically based on the pixel-wise error distribution, and propose novel score adjustment procedures for more precise gradient estimation. The source code is available at https://github.com/Bob-cheng/BadPart. Figure 1 presents our attack performance. In summary, our contributions are as follows:
@@ 2 Background and Related Work
Pixel-wise Regression Tasks. Pixel-wise regression tasks generate continuous values for each input pixel, differing from pixel-wise classification tasks like semantic segmentation (Wang et al., 2022; Liang et al., 2023, 2024; Liu et al., 2021), which assign a discrete class label to every pixel. Representative tasks of this type include monocular depth estimation (MDE) (Moon et al., 2019; Watson et al., 2019; Wang et al., 2023), optical flow estimation (OFE) (Teed & Deng, 2020; Ilg et al., 2017; Lu et al., 2023), and surface normal estimation (SNE) (Zeng et al., 2019; Lenssen et al., 2020; Bae et al., 2021). In MDE models, the output comprises the estimated pixel-wise distance between the 3D scenario and the camera capturing the input image, with each pixel corresponding to a distance estimation. OFE models use two consecutive image frames as input and output the estimated motion of pixels (the “optical flow”) between the two frames. For each pixel in the first frame, a 2D vector is estimated, indicating its offset to the corresponding pixel in the second frame. SNE models output the estimated orientations of the surfaces in the input image, described with the ”normal vectors” in 3D space that are perpendicular to the surface at the locations of pixels. Given that the first two tasks (MDE and OFE) have broader and more security-critical applications, such as autonomous driving (Karpathy, 2020), visual SLAM (Wimbauer et al., 2021), video composition (Liew et al., 2023), and augmented reality (Bang et al., 2017), our discussion primarily focuses on these two tasks. However, our proposed attack is a unified approach and can be readily applied to other pixel-wise regression tasks like SNE.
@@ 3 Problem Formalization
$\mathcal{F}(\mathbf{D})[i,j]$ $=Mean(\{\mathbf{\mathbf{D}}[k,1,i,j]\mid k=1...n\})$ (8)
$\mathcal{F}(\mathbf{D})[i,j]$ $=Mean(\{\|\mathbf{D}[k,:,i,j]\|_{2}\mid k=1...n\})$ (9)
@@ 4.2 Probabilistic Square Sampling
The sampling algorithm is designed to enhance the probability of selecting locations within the patch region that are more vulnerable to adversarial perturbations. As indicated in (Cheng et al., 2022) and (Cheng et al., 2024), the identification of vulnerable areas on images is critical for improving attack performance. To ascertain these vulnerable areas in the beginning, we employ an initialization phase ($K$ iterations), wherein we randomly sample the square location in a uniform manner. This period is denoted as $\mathbf{p^{(1)}}$ in Figure 2. After the initialization phase, the pixel-wise error map $\mathbf{M}$ caused by the latest patch is utilized as an indication of the vulnerable areas and we leverage this map to calculate the probability distribution of location sampling. Those areas with larger errors obtain higher probability. The background color of the patch $\mathbf{p^{(1)}},\mathbf{p^{(2)}}$ and $\mathbf{p^{(3)}}$ in Figure 2 illustrate the sampling probability distribution of square locations.
@@ 4.3 Score-based Gradient Estimation
15:     $[\Delta]^{+}\leftarrow[\Delta]^{+}/\operatorname{max}([\Delta]^{+})$; $[\Delta]^{-}\leftarrow[\Delta]^{-}/|\operatorname{min}([\Delta]^{-})|$
@@ 57 $\times$ 57 Patch (1%) 80 $\times$ 80 Patch (2%) 100 $\times$ 100 Patch (3%)
Models GA HB P-RS Ours WB GA HB P-RS Ours WB GA HB P-RS Ours WB
@@
  • * Bold texts denote the best attack performance among black-box methods. Abbreviations. GA: GenAttack (Alzantot et al., 2019), HB: HardBeat (Tao et al., 2023), P-RS: Patch-RS (Croce et al., 2022), WB: White-box Attack for reference.
5 Evaluation
In this section, we evaluate BadPart on 2 kinds of tasks including 7 subject models. We compare with 3 baseline black-box attack methods and a white-box one. A set of ablation studies are discussed and the source code is provided here: https://github.com/Bob-cheng/BadPart.
5.1 Experiment Setup
@@
Tasks & Models. We evaluate BadPart on two pixel-wise regression tasks of MDE and OFE. For MDE, we use Monodepth2 (Godard et al., 2019), DepthHints (Watson et al., 2019), SQLDepth (Wang et al., 2024) and PlaneDepth (Wang et al., 2023) as the target models. For OFE, we attack FlowNetC (Dosovitskiy et al., 2015), FlowNet2 (Ilg et al., 2017) and PWC-Net (Sun et al., 2018). These models are selected since they cover both the popular and SOTA models, and we attack the publicly available models with the highest input resolutions from their official repositories. MDE models were trained on the KITTI depth prediction dataset (Uhrig et al., 2017) and OFE models were trained on the KITTI flow 2015 (Menze et al., 2015).
@@
Baselines & Metrics. There are no direct baselines available due to a lack of previous research on black-box attacks against pixel-wise regression tasks. Hence we adapt three black-box patch attacks on image classification to our scenario as baselines. These include Patch-RS (Croce et al., 2022), a SOTA soft-label attack that employs random search and is akin to the patch-attack variant of SquareAttack (Andriushchenko et al., 2020); HardBeat (Tao et al., 2023), a SOTA hard-label attack that estimates gradients; and GenAttack (Alzantot et al., 2019), a conventional approach using genetic algorithms. We also compare BadPart with the white-box attack as reference. Discussion on more prior attacks can be found in Appendix D. We employ the mean depth estimation error (DEE, unit: meters) for MDE models and the mean end point error (EPE, unit: pixels) for OFE models as metrics to denote attack performance (the higher the better), which are aligned with the metrics used in prior white-box attacks on the two tasks (Cheng et al., 2022; Ranjan et al., 2019). They are calculated on the patched area and reflect the error caused by the generated adversarial noise.
@@ 5.2 Attack Performance
We compare the attack performance of BadPart with other baseline methods. We report the maximum attack performance of each method until convergence or after 1000K queries, whichever first. We also include the performance of white-box attacks as references. Three different patch sizes are evaluated and the patch locates at the center of the image. Table 1 reports the result. The first column denotes various pixel-wise regression models under attack, and the following columns represent attack performance of different methods and the white-box reference. As shown, BadPart obtains the best attack performance on all models under various patch sizes. The performance of BadPart is even close to the white-box attack reference on some models (e.g., Monodepth2, DepthHints and SQLDepth). Patch-RS has the second best attack performance while GenAttack performs the worse and has nearly no effect. Figure 3 presents qualitative examples of the attack performance of different methods on PlaneDepth and FlowNetC. As shown, the first column denotes the benign scene and the MDE and OFE output. The following columns show the $2\%$ adversarial patches generated by BadPart and baseline methods, as well as the model output. In the first row, when the patch generated by BadPart is applied, the depth estimation of the patched area is significantly further than the actual distance (darker color denotes further distance estimation). In comparison, the patches generated by other methods cause less impact. In the second row, patches generated by BadPart and Patch-RS have degraded the OFE performance significantly, making the result unusable. More qualitative and quantitative results are in Appendix C.
@@ 5.5 Attack Online Service
To evaluate the feasibility of BadPart in real-world applications, we conduct attacks on the Google API for portrait depth estimation (Google3DPortrait, ). Note that the model is not deployed by authors and we only query it for depth estimation. We utilize ten $256\times 192$ portrait images as training set and generate a $31\times 31$ patch using 50K queries to the API. Our attack goal is to minimize the estimated distance of the patched area on the portrait. BadPart has successfully reduced the mean depth estimation of the patched area by 43.5% from 0.431 cm to 0.243 cm. Figure 7 shows qualitative results. This adversarial example is available in the code repository, which can be uploaded to (Google3DPortrait, ) for efficacy validation. The adversarial patch can also be employed for beneficial purposes, such as privacy protection. By attaching the patch to personal images before publishing, individuals can prevent unauthorized use of their photos in such video composition services.
@@
Table 4: Universal attack performance with lower query budgets (2% patch size).
Query Monodepth2 (DEE: Meters) FlowNet2 (EPE: Pixels)
1K 1.923 2.050
5K 2.034 2.265
10K 2.392 2.272
20K 5.518 2.621
30K 12.299 2.851
Table 5: Single-image Attack performance with lower query budgets (2% patch size).
Query Monodepth2 (DEE: Meters) FlowNet2 (EPE: Pixels)
1K 3.200 2.063
5K 6.101 7.548
10K 9.060 13.141
20K 13.295 30.102
30K 24.235 41.562
5.6 Defensive Discussion
As pioneers in the exploration of black-box adversarial patch attacks against pixel-wise regression tasks, we find ourselves in uncharted territory with no existing defense techniques specifically tailored for this context. Nevertheless, there are defense strategies designed against black-box attacks on classification models. For example, certifiable defenses such as PatchCleanser (Xiang et al., 2022) employ a mask to traverse all input positions, monitoring output class mutations to identify the most suspicious position. Furthermore, universal adversarial patch detection methods (e.g., SentiNet (Chou et al., 2020)) depend on the input features responsible for the predicted class to locate the patch. However, their reliance on class output renders them unsuitable for direct application to the pixel-wise regression tasks. In contrast, query-based defense techniques, designed to detect malicious queries by black-box attacks, may be more applicable to our context. Blacklight (Li et al., 2022), a leading defense of this type, leverages the similarity among different query inputs to detect black-box attacks. Its primary strategy involves calculating the hash representation of each incoming query and identifying an adversarial query if the hash matches any previous one. Blacklight’s efficacy is contingent on the similarity between two images in consecutive queries, which is a major feature of single-image black-box attacks (e.g., SquareAttack (Andriushchenko et al., 2020)). However, BadPart is a universal adversarial patch attack that does not depend on sample similarity, and the randomness in different samples could potentially enhance the universal effectiveness of the generated patch. Hence we add random noise on each attack sample to by-pass the defense. Additionally, the high resolution of our input images further diminishes the efficiency and efficacy of such a defense. To assess the defensive performance of Blacklight on BadPart, we have incorporated it in our framework and evaluated its detection rate for varying amounts of queries. The results are presented in Table 3. As shown, for both MDE and OFE tasks, the detection rate remains zero under 800K queries, while the attack performance is not affected and continues to increase with more queries. We discuss more prior defenses in Appendix D, and the limitations and our future work in Appendix F.
5.7 Discussion on Lower Query Budget
Our attack requires 50K queries to cause an average depth estimation error (DEE) of 19.43 meters on Monodepth2, which employs a U-Net architecture with two ResNet18s. This is comparable to the SOTA universal black-box attack (HardBeat (Tao et al., 2023)), which requires 50K queries to achieve over 80% attack success rate on CIFAR-10 using a ResNet18 classifier. Other patch attacks requiring a smaller number of queries are for classification models, not pixel-wise regression models we are targeting, and they are not universal attacks. In particular, our scenario is characterized by the absence of class labels, precluding our ability to employ an image from the target class (e.g., an elephant) as a form of robust prior knowledge to initialize the patch on a source image (e.g., a clock), as did in DevoPatch (Chen et al., 2023). Additionally, we argue the one-time cost of 50K queries is affordable because it could be completed within 14 hours on the platform with a rate limit of 1 query per second (e.g., ClipDrop (Clipdrop, )), and even faster on Google 3DPortrait API (Google3DPortrait, ) without rate limitations. Actually, it only takes us less than 4 hours to attack this real-world API with 50K queries.
Please note that our method is an universal attack (rather than single-image attack), which is more challenging and practical. Although the patch generation process could require more queries, it is a one-time effort and the generated patch can attack arbitrary unseen images without further queries. As stated in the Attack Settings of Section 5.1, the attack performance of patches in our experiments is evaluated on an unseen test set, and Figure 11 shows the qualitative results, demonstrating the universal effectiveness of our patches across various scenes.
It is also worth mentioning that, unlike the classification tasks in which the attacks are either successful (i.e., output a wrong label) or unsuccessful (i.e., output the correct label), for regression models, the error caused by the attack (i.e., attack performance) increases continuously with the query times. That means, lower query budgets can still cause some amount of error. In table 4, we report the attack performance of our approach under lower query budgets. It is the same universal attack setting as our main experiments. As shown, using a lower query budget (e.g., 30K) could still cause noticeable errors on certain target models.
If the attacker aims on a single-image attack, he can easily adapt our method to such a scenario to reduce query times (just make the training set and the validation set only contain the target image). We conduct additional experiments to evaluate the attack performance of BadPart on a single image with different query budgets. Results are shown in Table 5. As shown, for both the MDE and OFE tasks, the errors caused by our single-image attack are already significant at 10K queries.